Problem 3 jan T & Find is if 1. = 8/450 V i a nr.
Answer #1

2.2 V -j2-2 M given Vo=8<45° v By voltage division principie, No= v1 ( 222 2 ; ) Vo 2 2-02 Vi = 812 Lov. Applying KCL at node- Vi (***) VG = 2 (873)(1104-1843) Ns = 17.884-18.45 V.